A firm evaluates all of its projects by using the NPV decision rule.

  

Year                 Cash Flow
0 –$29,000      
1 21,000      
2 15,000      
3 4,000      

  

a. At a required return of 17 percent, what is the NPV for this project?

  

b. At a required return of 33 percent, what is the NPV for this project?
Answer #1

a.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=21000/1.17+15000/1.17^2+4000/1.17^3

=31403.90

NPV=Present value of inflows-Present value of outflows

=31403.90-29000

=$2403.90(Approx).

b.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=21000/1.33+15000/1.33^2+4000/1.33^3

=25969.54

NPV=Present value of inflows-Present value of outflows

=25969.54-29000

=$-3030.46(Approx).(Negative).